The positive difference between k and 1/8 is the same as the positive difference between 1/2 and 1/3 . which of the following could be the value of k ?
a 1/7 b 1/6 c 7/24 d 23/24 e 4/3

Answer by mananth(16949)

k-1/8 = 1/2 -1/3
k= 1/2 -1/3 + 1/8
k=12-8+3 / 24
k=7/24